### v2

# SoK: Payment Channel Networks

Abstract
Payment Channel Networks (PCNs) have been proposed as an alternative solution to the scalability, throughput, and cost overhead associated with on-chain transactions. By facilitating offchain execution of transactions, PCNs significantly reduce the burden on the blockchain, leading to faster transaction processing, reduced transaction fees, and enhanced privacy. Despite these advantages, the current research in PCNs presents a variety of research challenges that require further exploration.
In this paper, we survey the recent work in several aspects of PCNs, such as pathfinding and routing, virtual channels, state channels, payment channel hubs and rebalancing. This survey aims to provide the reader with a detailed understanding of the current state-of-the-art in PCN research, highlighting a few important advancements. Additionally, we highlight the various unresolved issues in the area of PCN research.
Specifically, this paper seeks to answer the following crucial question: What are the various interesting and non-trivial challenges in PCN research that require immediate attention from the academic and research community? By addressing this question, we aim to identify the most pressing problems and future research directions that interested readers can immediately work on. Through this analysis, we hope to inspire researchers and practitioners to tackle these challenges to make PCNs more secure and versatile.
Index Terms:
Payment channel network, Blockchain, Layer-2
I Introduction
(a) Payment channel opening
(b) Alice sending 30 coins
(c) Local balances get updated
(d) Posting final balances to blockchain
Figure 1: Payment channel operations
Cryptocurrencies and cryptocurrency based transactions have become increasing popular. Currently, the total market value of all cryptocurrencies in use has surpassed 2.5 Trillion USD. The cyrptocurrency market is increasing at a rate of $\approx$ 8.00% every year [1]. This rise in popularity can be attributed to the following reasons: 1) cryptocurrency transactions can be carried without the presence of a trusted entity. Fiat currency based transactions on the other hand, require the presence of a trusted financial organization such as a bank. 2) cryptocurrency transactions do not subject the user to any limits on the number and type of transactions. Fiat currency based transactions are limited in their amount and number, depending on several factors such as the currency, geographical location, etc. [2].
Each transaction posted to the Bitcoin blockchain takes around 7 seconds to be validated [3, 4]. The procedure of validation involves verifying that the transaction posted to the blockchain contains all the required fields and the signature of the user creating the transaction tuple is valid. Once the validation procedure is successfully completed, the transaction is included in a block that would be mined on the blockchain at some time in the future. The process of mining the block successfully takes $\approx$ 2 hours [5] (as of June $13^{th}$ 2024). This delay in the transactions and blocks getting confirmed is termed as the blockchain scalability problem [6]. In contrast, Visa, a company which globally processes transactions using fiat currencies, processes around 24,000 transactions per second [7]. Due to the delay in transaction processing caused by the blockchain scalability problem, blockchain-based transactions cannot process payments instantaneously.
As an alternative to processing transactions by posting to the blockchain, payment channels have been proposed. Two parties with the intent of processing payments between them open a payment channel by creating a transaction tuple called the funding transaction. This funding transaction contains the initial deposits from both the parties. These initial deposits are also called as the initial balances of the parties in the payment channel. The sum aggregate of these initial balances is called as the channel capacity. The funding transaction contains the signatures of both the parties involved in the payment channel making it a 2-2 multi signature transaction. This means that the funds in the funding transaction cannot be spent without the signatures of both the parties. This funding transaction is validated and included in a block. Once this block has been successfully mined and confirmed on the blockchain, the payment channel is opened between the two parties. The two parties can now be involved in an unlimited number of transactions with each other as long as the amount of a single transaction does not exceed their local balances.
An example of a payment channel is given in Figure 1(a). Two users (also called nodes) Alice and Celia deposit 50 coins each into a 2-2 multi signature transaction. This transaction is posted to the blockchain, upon which it is validated and included in a block. The block is mined and confirmed, at which time a payment channel is said to open between Alice and Celia. The sum aggregate of the individual balances of Alice and Celia, which is the channel capacity, is 100 coins. Alice making a payment of 20 coins to Celia is shown in the Figure 1(b) and the updated balances of Alice and Celia are shown in Figure 1(c). After the payment has been made, the channel capacity between Alice and Celia still remains constant at 100 coins. In this manner, Alice and Celia can be involved in an unlimited number of payments between each other. Each payment made in the payment channel in an off-chain manner contains the signatures of both Alice and Celia. When either Alice or Celia decide to close the payment channel, they post their final balances to the blockchain and the payment channel is closed as shown in Figure 1(d).
For each off-chain transaction in the payment channel, both the parties involved in the payment channel create a commitment. This commitment is essentially an agreement for the new balances signed by both the parties. Exchanging of commitments signifies that both the parties have agreed to the change in their respective balances. Each pair of commitments (for each transaction), contains a unique sequence number called the revocation sequence maturity number. For each new transaction made in the payment channel, the sequence number of the prior transaction is invalidated by revocation keys of both the parties. These revocation keys are created by both the parties before opening of the payment channel. If a malicious party in the payment channel broadcasts an older balance to the blockchain, the honest party in the channel has a certain time period during which it can contest this behavior on the blockchain. Before this time period expires, the honest party in the payment channel will broadcast the revocation of this old state signed by both the parties. The broadcasting of this revocation to the blockchain prevents the malicious party from stealing funds of the honest party.
The idea of a payment channel that exists between two parties can be extended to a number of nodes, creating a network of payment channels, called a payment channel networks or PCN. The advantage of PCNs is, they enable users that are not connected by a direct payment channel to make payments between each other in an off-chain manner. An example PCN is shown in Figure 2. In the figure, consider Alice who intends to make a payment to Hector, with whom she does not share a payment channel. The naive way to process this transaction would for Alice to open a payment channel with Hector, which involves Alice making an expensive blockchain write for the channel opening. Each payment channel opening costs 2.4 USD for blockchain writes [8, 9, 10]. If Alice intends to send an amount of 1 coin to Hector, it may not be economical for her to open a direct payment channel. Alice can make use of the PCN and make a payment to Hector by forwarding the payment along the path Alice $\rightarrow$ Celia $\rightarrow$ Michael $\rightarrow$ Rajiv $\rightarrow$ Charlie $\rightarrow$ Garcia $\rightarrow$ Hector. This process of using intermediate nodes in a PCN to forward to the payment to the intended destination is called as routing in payment channel networks.
Figure 2: Payment channel network
Motivation and timeliness of PCN research: A significant advantage of PCNs is their capability to facilitate micro-payment transactions, with minimum amounts as low as $10^{-7}$ BTC [11]. In contrast, the average transaction cost for a single on-chain transaction on the BTC blockchain is approximately $4.612$ USD as of June 2024 [12]  [^1] Opening a payment channel requires two inputs (for funds from both the parties) and one output (the funds are locked in a single 2-2 multi signature address). Where as making a payment by posting it to the blockchain requires one input (from the party that intends to make the payment) and two outputs (one to the receiver and one to the validator/miner for his fees/reward). This difference makes the transaction cost associated with opening a payment channel cheaper than making a payment using a blockchain write., regardless of the transaction amount. This cost can be avoided by using the off-chain PCNs, which incurs no additional fees. Additionally, transactions on the BTC blockchain take around 2 hours to be confirmed as of June 13, 2024 [5], whereas PCNs can process transactions instantly. An example of a real-world PCN is the Lightning Network ($\mathsf{LN}$) on the Bitcoin blockchain [13], which has a 24-hour trading volume of \$63,200 and a market capitalization of \$7 million [14] as of June 2024. These figures reflect $\mathsf{LN}$’s size and growth.
Contributions:
1) We qualitatively compare the recent work in several aspects of PCNs, viz. pathfinding and routing, virtual channels, state channels, payment channel hubs and rebalancing using several relevant properties (metrics) along with providing a reasoning why these metrics have been chosen for comparison.
2) In this paper, we point out the open problems in all the areas that we survey and we also discuss why solving those problems is a hard research challenge.
Outline: In Section II, we start with describing the concept of pathfinding and routing in PCNs and qualitatively compare work published in that area. In Section III, we describe virtual channels which have been proposed to address issues with multi-hop routing in PCNs, and compare work in this area. In Section III, we describe and compare state channel protocols, which are a more generalized adaptation of virtual channels and can facilitate the execution of arbitrary applications between nodes in the PCN (not just payments). In Section V, we cover payment channel hubs, which are similar to virtual channels and state channels, but facilitate a different use case for payments in PCNs. All of the aforementioned PCN mechanisms consume one common resource: the local balance of a node in a PCN. In Section IV, we discuss rebalancing, which addresses the important function of replenishing channel funds in the PCN. In Section VII, we present the reader with the current research gaps in all of the aforementioned areas and in Section IX, we conclude the paper.
Prior work: Prior works by Khojasteh et al. [15] and Erdin et al. [16] survey the work done only in the area of pathfinding and routing protocols and their privacy aspects in PCNs. Whereas, in this paper we cover the entire spectrum of PCN research: rebalancing, virtual channels and state channels, pathfinding and routing. Apart from this [15, 16] do not provide any information about the open problems in PCNs, which we do in our work. The SoK by Gudgeon et al. [17], surveys several layer-2 protocols, whereas, we focus exclusively on PCNs. Furthermore [17] was published in 2020 and does not cover most of the recent work published in PCNs.
TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs.
Protocols Concurrency Privacy Topology privacy Avoids source routing Decentralized Atomicity Disjoint graphs Fees
MPCN-RP [18] $\checkmark$ $\checkmark$ (flat)
Eckey et al [19] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Vein [20] $\checkmark$ (dynamic)
Auto tune [21] $\checkmark$ (flat)
Kadry et al. [22]
FSTR [23]
SilentWhispers [24] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Blanc [25] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SpeedyMurmurs [26] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Spider [27]
Flash [28] $\checkmark$
Coinexpress [29] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Webflow [30] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Robustpay [31] $\checkmark$ $\checkmark$ (flat)
Robustpay+ [32] $\checkmark$ $\checkmark$ (flat)
RACED [33] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Cheapay [34] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (flat)
Auroch [35] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (dynamic)
Yang et al. [36] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SPRITE [37] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
II Pathfinding and Routing
Motivation: One of the areas in PCNs which has garnered significant attention from the academic community is pathfinding and routing. It is defined as the process of finding a path comprising of several nodes from a sender to a receiver in a PCN along which the payment intended to be sent by the sender can be forwarded to the receiver and forwarding the payment along that path.
Intuitively, it may seem that well-known pathfinding and routing protocols from the networks domain can be easily applied to PCNs. Unfortunately, there are several problems with this.
Even though PCNs are decentralized and distributed peer-peer networks, pathfinding and routing protocols of traditional networks cannot be applied to solve the pathfinding and routing problem in PCNs, because of fundamental differences between intent and the network features: 1) Transfer of data and transfer of money: The intent behind traditional network’s pathfinding and routing protocols is to transfer data between connected devices. In PCNs, the intent to transfer money between users in a decentralized and distributed manner. 2) Change in bandwidth and change in monetary state: When data is transferred between connected devices using network pathfinding and routing protocols, the bandwidth, which is represented by the link weights between the nodes does not change. However, transfer of money between users in a payment channel permanently alters the state of the sender and receiver involved by subtracting and adding money from and to their link weights respectively. 3) Latency and routing fees: The cost of sending data in network pathfinding and routing protocols is measured in terms of latency, which is defined as the time delay experienced in transferring data from one device to another. The delay in data transfer is experienced by the receiver and the cost of the delay is usually paid by the sender. The cost of sending payments in a PCN is measured in terms of money, called as routing fees. This encourages greedy behavior and makes the PCN prone to several attack vectors. Due to these reasons, pathfinding and routing in PCNs is a fundamentally different and a hard research challenge.
The properties in Table I represent the fundamental principles of fiat currency transactions and on-chain payment mechanisms, which we want reflected in off-chain payments. These properties are generally agreed upon in the literature by several works such as [33, 35, 37, 25, 26, 24] as common evaluation metrics for pathfinding and routing protocols in PCNs. Fulfilling these properties while providing efficient pathfinding and routing is a non-trivial challenge, and necessitates the design of new pathfinding and routing protocols. Several elegant pathfinding and routing protocols have been proposed in the literature. In Table I, we present a qualitative comparison of these routing protocols with respect to the properties they achieve.
Ideal properties:
1) Concurrency: Concurrency is achieved when a pathfinding and a routing protocol enables the nodes to forward more than one payment simultaneously.
2) Privacy: Privacy is maintained when a node’s identity is known only to its immediate neighbors and not to the entire network.
3) Topology Privacy: Topology privacy is preserved when no single node has knowledge of the entire network topology.
4) Avoids Source Routing: Source routing is avoided when the sender does not determine the path to the receiver.
5) Decentralization: Decentralization is achieved when there are no centralized or trusted entities responsible for constructing paths for senders.
6) Atomicity: Atomicity is ensured when the payment is routed all the way from the sender to the receiver or the payment is not routed at all.
7) Disjoint Graph Applicability: A pathfinding and routing protocol is considered applicable to disjoint graphs if it functions even when the network graph consists of islands.
8) Fees: Routing fees is the amount a node charges for forwarding the payment to the next node along a path from the sender to receiver. This fees can be charged in two ways. Flat/fixed fees means that the fees charged for routing payments remains the same irrespective of the transaction amount being routed. If the fees charged by a node varies according to the transaction amount, it is referred to as dynamic fees, typically a percentage of the amount being transacted.
III Virtual channels
(a) Virtual channel example [38, 39]
(b) Multi-hop virtual channel (the dots between Jill and Alice depict several intermediate nodes)
(c) Recursive virtual channel
Figure 3: Different types of virtual channels
TABLE II: Comparison of Virtual Channel Protocols
Protocol BC Validity Fee Privacy Off-chain dispute resolution Recursive Multihop
Perun [38] TC Limited
Jourenko et al. [40] UTXO Limited $\checkmark$
Aumayr et al. [41] UTXO Limited $\checkmark$ (fixed)
Elmo [42] UTXO Unlimited $\checkmark$
Donner [43] UTXO Unlimited $\checkmark$ (time based) $\checkmark$ $\checkmark$
Jia et al. [44] UTXO and TC Limited $\checkmark$ (fixed)
Eckey et al. [45] TC Limited $\checkmark$
Generalized state channels [46] TC Limited $\checkmark$
Transactions in PCNs are routed from the sender to the receiver using a path of intermediate nodes. This pathfinding and routing mechanism requires the nodes along the payment path to be available for a transaction to be processed. Nodes, however, sometimes can choose to go offline or there can be network/service disconnections. This might cause the transaction to fail. Virtual channels, which are built on top of existing payment channels, provide a solution to this problem. Virtual channels facilitate payments between a pair of nodes using an intermediate node. The intermediate node has individual payment channels open with the other two nodes. The intermediary and the pair of nodes lock coins with each other in their respective payment channels to establish a virtual channel. Once the virtual channel has been established, the two nodes can be involved in a unlimited number of payments. These payments can be processed even if the intermediate node goes offline.
The first virtual channel construction, Perun, was proposed by Dziembowski et al. in [38]. Assume there are two users, Alice and Bob, along with an intermediary as depicted in Figure 3(a). Alice establishes a payment channel, $\alpha_{A}$, with the intermediary by locking $Y_{A}$ coins as collateral; the intermediary contributes $Y_{I}$ coins into this channel. Similarly, Bob locks $Z_{B}$ coins into his channel, $\alpha_{B}$, with the intermediary, which also deposits $Z_{I}$ coins from its possession. This marks the completion of the opening phase, effectively locking the coins belonging to Alice and Bob. Following this, Alice and Bob create a virtual channel to facilitate transactions between them. The virtual channel will include coins deducted from both Alice’s and Bob’s payment channels. Specifically, $X_{A}$ coins will be subtracted from $Y_{A}$ on Alice’s side and deposited into the virtual channel, while $X_{B}$ coins will be subtracted from $Z_{B}$ and will be deposited from Bob’s side. Now Alice and Bob can be involved in an unlimited number of payments without the intermediary being online.
The idea of a virtual channel between a pair of nodes involving a single intermediary/intermediate node has been extended to establishing a virtual channel recursively over several hops involving several intermediaries, leading to the construction of a recursive virtual channel [40, 42, 43]. An example of a recursive channel is shown in the Figure 3(c). In the figure Trent and Alice are a pair of nodes who intend to establish a virtual channel between them. They do in three steps as follows; 1) Trent and Hu construct a virtual channel $V_{1}$ with Rajiv as an intermediary. 2) Hu and Alice construct a virtual channel $V_{2}$ with Jill as an intermediary. 3) Trent and Alice establish a virtual channel between them $V_{3}$.
A multihop virtual channel is constructed by establishing a single virtual channel between a pair of users over a path comprising of several intermediate nodes. An example of multihop virtual channel is shown in the Figure 3(b). Rajiv and Alice establish a single virtual channel $V$ between them.
Motivation: We have already described how PCNs offer a solution to the blockchain scalability problem by facilitating instantaneous payments in an off-chain manner. PCNs however, have few limitations: 1) The path generated in a PCN to process a payment between a sender and a receiver usually contains one or several intermediate nodes. All the nodes along the path are required to be available to route the payment until the receiver. This may not be possible at all times since, nodes can be disconnected from the network due to various reasons such as power outages, DoS attacks, etc. In a virtual channel, once the coins of the intermediary(ies) are locked, it can choose to remain offline without affecting the transaction processing. 2) Each node along the path from the sender to the receiver charges a fees called as routing fees/forwarding fees for forwarding the payment to the next node along the path. The sender has to pay all the nodes their respective forwarding fee along the path. This increases the overall cost of routing the payment. Virtual channels provide a solution to this problem, since they do not charge fees on a per transaction basis. Instead, the fees is charged based on a time epoch basis (the time for which the coins of the intermediary/intermediaries are locked in the virtual channel) or a fixed fee can be agreed in advance by all the parties involved in the virtual channel construction. 3) The time taken to route a payment from a sender to a receiver is directly proportional to the path length. A longer path would result in an increase in the overall pathfinding and routing time. Virtual channels do not process payments through a path comprising of several intermediate nodes, hence the overall transactions’ pathfinding and routing time does not depend on the number of intermediate nodes between the sender and the receiver.
Virtual channels should not add an unnecessary burden to users, and should mirror the operations of payment channels as closely as possible, with the added benefit of no on-chain transactions at all, while maintaining comparable security/privacy properties. We now give the properties desired from an ideal virtual channel, and compare the works in this area in Table II on the extent to which they achieve these properties.
Ideal properties:
1) BC: BC is the blockchain on which the corresponding virtual channel protocol can be deployed. TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO based scripting mechanism such as Bitcoin. Ideally, a virtual channel protocol should be deployable on any blockchain without the need for any additional programming/scripting requirements. The virtual channel protocol should not require any modifications to the existing consensus mechanisms of the underlying blockchains.
2) Validity: This determines the validity of the virutal channel. A limited validity means that the virtual channel is valid for a predetermined time period (which is decided by parties involved in the virtual channel). Unlimited validity means that the virtual channel can stay open until the parties involved decide to initiate closing. This provides better and efficient transaction processing between users who intend to have frequent payments made between each other.
3) Fee: This metric determines if the virtual channel takes into account the fees charged by the intermediate node(s) involved in the channel’s establishment. The fixed fee model implies that a predetermined, fixed fee is paid to the intermediate node(s) which is agreed upon by all the parties in the virtual channel and the time based fee model implies that a fixed fee is paid to the intermediary which depends on the time for which the virtual channel stays open. Having an well-defined fee structure will motivate the intermediary/intermediaries to participate in virtual channel creation.
4) Privacy: Similar to payment channels, we want privacy of nodes for virtual channels, which is said to be achieved when the identities of the nodes in the virtual channel protocols are not known to anyone except for their immediate neighbors.
5) Offchain dispute resolution: This metric determines if the transaction disputes in a virtual channel require a blockchain write.
6) Recursive: A virtual channel is said to be recursive or hierarchical if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by recursively constructing virtual channels on top of existing virtual channels from sender to receiver. We give an example of a recursive virtual channel in Figure 3(c). 7) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channel between any pair of intermediate nodes along the path. We give an example of a multihop virtual channel in Figure 3(b).
(a) State channel opening
(b) Chess game using state channel
(c) State channel close
Figure 4: State channel example with a chess game decentralized application (DApp)
TABLE III: Comparison of State Channel Protocols
Protocol BC General/purpose state channels Security Graceful exit Off-chain dispute resolution
Pisa [47] TC $\checkmark$
Hydra [48] UTXO $\checkmark$ $\checkmark$ $\checkmark$
Sprites [49] UTXO $\checkmark$ $\checkmark$ $\checkmark$
Aumayr et al. [50] UTXO, TC $\checkmark$ $\checkmark$
Origami [51] TC $\checkmark$ $\checkmark$
Forcemove [52] TC $\checkmark$
IV State channels
A state channel is an off-chain protocol that can facilitate execution of an arbitrary decentralized application, also called as DApp (such as a two player game) between two users in a decentralized and distributed network. An example of a state channel is given in Figure 4(a). Let us consider two users Alice and Celia who intend to play a game of chess. They initially interact with the chess game application deployed on the blockchain by a third party service provider. Both Alice and Celia deposit 30 coins each from their possession into a 2-2 multi signature transaction (step 1). Upon locking the funds, the chess game is instantiated between Alice and Celia by the DApp (step 2). Alice and Celia engage in a series of moves for the chess game (step 3, step 4) and let us assume Alice eventually wins (step 5) as shown in Figure 4(b). Alice gets paid 15 coins as the prize money in Figure 4(c). In this game, Alice and Celia countersign each other’s moves until there is a winner or a draw. In the event of malicious activity by one party (such as undoing a prior move), the honest party will broadcast all the moves to the blockchain for dispute resolution.
Motivation: State channels allow the already existing payment channels to facilitate execution of arbitrary applications such as games, e-commerce, etc. This leads to the PCN becoming more versatile. State channels should be deployable on any blockchain regardless of the underlying programming/scripting requirements and should be able to facilitate the execution of any decentralized application between the parties involved, while maintaining comparable security properties (in comparison to payment channels) and should ideally be able to resolve disputes without accessing the blockchain. We now give the properties desired from an ideal state channel, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of various state channel protocols in Table III. We describe the metrics used for comparison below.
Ideal properties:
1) Blockchain: This is the blockchain on which the corresponding state channel protocol can be deployed. TC is a blockchain that supports a Turing complete programming language and UTXO if a blockchain that supports a UTXO based scripting.
2) General purpose state channels: This is the ability of the state channel to facilitate the execution of any application in an off-chain manner by using a single application-agnostic instance of state channel on the blockchain.
3) Security: Security is achieved by a state channel protocol if no honest party loses funds because of malicious behavior by other parties in the protocol.
4) Graceful exit: A state channel protocol is said to employ graceful exit if it has clear and well-defined mechanisms for nodes joining or leaving the state channel.
5) Off-chain dispute resolution: This metric indicates if a dispute that occurs between the parties involved in the state channel can be resolved without accessing the blockchain.
V Tumblers
TABLE IV: Comparison of Tumblers
Protocol BC
Relationship
anonymity
Privacy against
aborts
Time epochs
Dynamic
corruption
Atomicity
Value
privacy
Variable
amount
BOLT [53] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Nocust [54] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MIXCT [55] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
A2L [56]
UTXO and
TC
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Blindhub [57]
UTXO and
TC
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Teechain [58] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Accio [59] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Turbo [60] TC $\checkmark$ $\checkmark$
Boros [61] TC $\checkmark$ $\checkmark$ $\checkmark$
Tumblebit [62] UTXO $\checkmark$ $\checkmark$ $\checkmark$
Motivation: A payment channel hub is a layer-2 mechanism that facilitates transaction processing between a sender/receiver pair that do not have a payment channel open between them. The intermediary which facilitates payments is called a tumbler. The tumbler has multiple payment channels open with several nodes in the network to facilitate payments. The hubs can potentially be useful in facilitating micropayments between a buyer and seller, without either of them needing to open a payment channel with each other.
Payment channel hubs can be classified into two types: on-chain and off-chain. Early hubs, also called hubs or mixers, were on-chain [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80], but they all suffered from scalability issues due to having to post each transaction on the blockchain. This led to the development of offchain payment channel hubs for specific blockchains, e.g., Bolt [53] for Zcash, Nocust [54] and MixCT [55] for Ethereum. The most general-purpose payment channel hubs are Tumblebit [62], $A^{2}L$ [56], and Blindhub [57].
A payment channel hub should be able to facilitate payments between a sender and receiver, who do not have a payment channel open between them such that the hub cannot link a given transaction amount to a particular sender/receiver pair. Furthermore, the payment channel hub should also guarantee the fundamental property of atomicity (ensuring that the payment is sent to the receiver or it does not go through at all). We now give the properties desired from an ideal payment channel hub, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of off-chain payment channel hubs in the Table IV. We describe the comparison metrics below:
Ideal properties:
1) Blockchain: BC is the blockchain on which the corresponding Tumbler protocol can be deployed. TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO based scripting mechanism such as Bitcoin.
2) Relationship anonymity: It ensures that the relationship between a sender and a receiver for a given transaction should not be known to the tumbler.
3) Privacy against aborts: Tumbler should not be able to deduce the identities of a sender/receiver in case of a transaction abort, regardless of which party is responsible for the abort.
4) Time epochs: The tumbler processes transactions in discrete fragments of time called epochs, i.e., transactions only take place during a time epoch. A payment channel hub that supports transaction processing based on time epochs is not efficient because nodes cannot process payments between time epochs, which can be of arbitrary length. The duration between time epochs depends on several factors such as the unit of the time epoch (absolute time or relative time measured in terms of block height), number of users that will most likely involved in the next time epoch, etc. 5) Dynamic corruption: The payment channel hub should be able to handle the deviation of any party from the protocol at any point during its execution. i,e the set of adversaries should not be static during the protocol execution.
6) Atomicity: Atomicity ensured when a payment is routed all the way from the sender to the receiver or the payment is routed at all.
7) Value privacy: Value privacy is ensured by a payment channel hub, when, given a transaction amount, the tumbler cannot link it to a sender/receiver pair.
8) Variable amount: The tumbler should be able to process transactions of any value.
TABLE V: Comparison of Rebalancing Protocols
Protocol BC Trusted entity Graph compatability Privacy
Revive [81] TC $\checkmark$ (C) Cycles only
Hide & Seek [82] UTXO $\checkmark$ (D) Cycles only $\checkmark$
Subramanian et al. [83] TC and UTXO Agnostic $\checkmark$
Cycle [84] TC Cycles only $\checkmark$
Rebal [85] UTXO Cycles only $\checkmark$
Shaduf [86] TC Agnostic $\checkmark$
Chen [87] et al. UTXO Cycles
Musketeer [88] UTXO Cycles
VI Rebalancing
Motivation:Rebalancing in payment channel networks involves redistributing the balance within an existing payment channel or across multiple channels to maintain or improve the network’s liquidity and usability. This is done to ensure that both parties in the channel have sufficient funds on their respective sides to send and receive payments. Rebalancing is crucial to payment channel networks. If the link weights of nodes become zero/depleted as a result of being involved in transactions, it will prevent them from taking part in further payments until those link weights are replenished. Such nodes are called as dormant nodes in the network. A PCN which has a large number of such dormant nodes experiences an overall reduction in transaction throughput and an overall reduction in liquidity. Hence, the link weights of the nodes in the PCN need to be rebalanced to prevent failure of transactions due to lack of liquidity. A rebalancing protocol should be be able to replenish the link weight of dormant nodes in the PCN irrespective of the PCNs topology, while not having to employ any trusted entity and should not compromise privacy of nodes. We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of rebalancing protocols in Table V. We describe the metrics as follows:
Ideal properties:
1) Blockchain: This is the blockchain on which the corresponding rebalancing protocol can be deployed. As it can be inferred from the Table, the protocols can either be deployed on the Ethereum or the Bitcoin blockchains.
2) Trusted entity: This metric determines the nature of trust required among parties to deploy the rebalancing protocol. A centralized trust (C) indicates the presence of a single trusted entity and distributed trust (D) indicates the distribution of trust across various entities.
3) Graph compatibility: This indicates the nature of the topology of the network on which a rebalancing protocol can be deployed. Cycles indicate that the rebalancing protocol can only deployed on cyclic graph and agnostic indicates that the rebalancing protocol can be deployed irrespective of the network topology.
4) Privacy: Privacy is said to be achieved when sensitive information such as the local balance and the identities of nodes are not known to anyone except for their immediate neighbors.
VII Research Gaps & Open Problems
In this section, we highlight the gaps in research published until now in the areas of pathfinding and routing, virtual channel construction, payment channel hubs, and state channels. This section provides an overview of the current research problems in these areas along with a discussion on why is developing a solution to these problems an interesting research challenge.
VII-A Gaps in pathfinding and routing for PCNs
We now identify the gaps in research in pathfinding and routing protocols for payment channel networks.
1) Super node liquidity validation in PCNs: A super node is a highly connected node (it has payment channels open with numerous other nodes in the network) and also has a high liquidity in those channels. These super nodes aid other nodes in pathfinding and routing of payments. These super nodes are called as a trampoline nodes in $\mathsf{LN}$ [89, 90, 91, 92]. Several other works in the literature refer to super nodes as routing node, routing helper, landmark node, routers, etc [25, 24, 27, 37, 33]. One of the main problems with the super nodes is that the sender has no way to know whether the super node possess the local balance (liquidity) to route the payment of the sender. The local balance of a super node or any node in a PCN is a private value and should not be known to any node except for its immediate neighbor. Currently, if a super node does not have enough liquidity to route the payment of a sender, the transaction fails and it has to be retried by the sender. In $\mathsf{LN}$, one of the most widely used PCN, this is a significant problem. Sometimes the sender might have to keep retrying for $\approx$ 1 hour to have a successful payment [93]. The main goal of PCNs is to facilitate instantaneous payments and these transaction retries render such payments almost impossible. It might greatly benefit the sender if the it has a mechanism to validate whether a super node has enough liquidity (balance) to route its payment without violating any privacy concerns in the network. Achieving this is hard because the local balance a node has with its immediate neighbors in a PCN is considered private information and the sender has to know whether the super node possess enough liquidity without comprising privacy.
2) Channel verification in a PCN: To be a part of any PCN, nodes will open payment channels with other nodes in the network. Two nodes open a payment channel between them by posting a transaction to the blockchain. This transaction can be posted on the blockchain (as in UTXO based blockchains like BTC) or this transaction can be posted as a function call to an existing smart contract (as in blockchains that support smart contracts like ETH). In the most popular PCN, $\mathsf{LN}$, the channel verification procedure is very inefficient. It involves the node who intends to verify a channel to request the block in which the channel opening transaction has been included, verifying whether the transaction has been successfully executed by the validator/miner and finally verifying if the channel opening transaction corresponds to a 2-2 multi signature address on the blockchain. The verifier performing these steps is inefficient since all these steps will have already been performed by the miner. According to the current verification process, any node would need to perform all of these steps for each payment channel it has open in the network. Consider the following scenario: a node $j$ in the PCN has payment channels open with $m$ nodes in the PCN. Furthermore, assume that each neighbor opens $k$ new channels for each time epoch. For this example, assume the value of the time epoch is 1 block. This would mean that the total number of channels the node $j$ has to verify a total of $m*k$ channels for every block. This requires the node $j$ to access the blockchain a total of $m*k$ times to perform channel verification. Finding a way to do this without blockchain access and in a blockchain agnostic manner in a hard research challenge.
VII-B Gaps in Tumblers
We will now describe the research gaps in tumblers. 1) Privacy against aborts: In a given a transaction, the tumbler should not be able to correctly identify the sender/receiver pair involved. Currently, there is only one construction [59] that guarantees this. However, this construction works only for blockchains that support Turing complete programming languages. Currently, there is no such construction for UTXO based blockchains that guarantees privacy against aborts. This is hard because if the sender aborts a transaction, the tumbler will know which receiver did not receive the payment. This is hard because, a tumbler mixes multiple transactions together to ensure unlikability among payments. Also, a tumbler should enforce atomicity so that the payment either goes through completely or does not go through at all. In the case of an abort, an adversary who controls the tumbler might be able to deduce which receiver amongst the ones involved did not receive a payment and will be able to link the corresponding transaction amount to that receiver.
2) Time epochs: A tumbler should be able to facilitate payments without having a predefined time interval or epoch. Currently, the majority of payment channel hub constructions process payments in discrete epochs to increase the size of their anonymity sets. An anonymity set refers to a set of users (senders/receivers) whose identity is indistinguishable from each other. The size of this set is proportional to the privacy of sender/receiver in the transaction. When the interval between two consecutive time epochs is sufficiently large, it will promote more users to take part in the transactions within a time epoch, which improves privacy. Not having time epochs results in smaller anonymity sets and hence could possibly weaken privacy. The trade off is between privacy and flexibility.
3) Dynamic corruption: A tumbler should ideally be resistant to dynamic corruption. This means the protocol should be resilient to parties deviating from their ideal execution during the beginning, progress, and ending of the protocol’s execution. Current tumbler constructions only provide support for static corruption, which may be unrealistic from a practical standpoint. Security against dynamic corruption is hard to achieve because the tumbler mixes several transactions together to preserve anonymity. Since the tumbler should not be able to link transaction amount to a sender/receiver pair (ideally), if the sender/receiver becomes malicious during the protocol’s execution, ideally, that transaction must be aborted to prevent financial loss to the honest party. However, the tumbler should not be able to link a transaction to a sender/receiver pair, making security against dynamic corruption a hard research problem.
VII-C Gaps in Virtual Channels
We will now describe the open problems in the area of virtual channels.
1) Blockchain agnostic: A virtual channel should ideally be blockchain agnostic, meaning it should be deployable regardless of the underlying scripting or programming language structure supported by the blockchain.
2) Well defined fee structure: The intermediary/intermediaries involved in the virtual channel construction additionally lock coins apart from the ones locked in the underlying payment channel. It is necessary to have a well-defined fee structure for this additional service. Currently, nodes get paid a routing fees/forwarding fees for every transaction they process. This fee is dependent on the transaction amount being processed. In the case of virtual channels, having a well-defined fee structure is difficult due to the following reasons: 1) The fee structure should take into account the amount of funds and the time for which these funds of the intermediary/intermediaries are locked in virtual channel. 2) It also needs to take into account the routing/forwarding fees an intermediary could have earned by not locking up the coins in the virtual channel. The fees for an intermediary should be ideally be more than the routing fees it would have earned in order to motivate nodes to be involved in virtual channel construction. Currently, no virtual channel construction has a well-defined fee structure.
3) Off-chain dispute resolution: Ideally a virtual channel protocol should provide the flexibility to resolve any disputes without accessing the blockchain. We have seen in Section I, that blockchain writes are expensive. Having an off-chain dispute resolution is a hard research challenge due to the following reasons; 1) Any form of layer-2 disputes are usually resolved on the blockchain. Since underlying consensus in blockchain usually guarantees resolution of a dispute raised within a specific time period. In an off-chain scenario, there is no such mechanism which guarantees this. None of the virtual channel constructions have a provision of off-chain dispute resolution.
4) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channel(s) between any pair of intermediate nodes along the path. Currently, no VC construction has this capability. This is hard because, multihop virtual channel involves designing a mechanism for sequentially locking coins across a path comprising of several intermediate nodes. Furthermore, the construction of a multihop virtual channel should also ensure that any transaction does not effect the overall balance neutrality (no node apart from the receiver along the path gains or loses any money). Also, the goals of privacy, atomicity, offchain dispute resolution need to be enforced.
VII-D Gaps in state channel protocols
We will now describe the gaps in the area of state channel protocols.
1) General purpose state channels: An ideal blockchain should be able to facilitate deployment and execution of several decentralized applications without requiring the users to access the blockchain for any purpose other than opening and closing the state channel. Currently, there are two such constructions [49, 48] which incorporate code reusability into their design. These two constructions support UTXO based blockchains and have no provision for offchain dispute resolution.
2) Offchain dispute resolution: Current state channel protocols use an onchain transaction or a function call to the onchain smart contract in the case of dispute resolution. An ideal state channel protocol should be able to facilitate dispute resolutions in an offchain manner. This is a hard research challenge since disputes on the blockchain are usually resolved by the underlying consensus mechanisms. In an offchain scenario, there is no such consensus that guarantees transaction validity. Also, in an offchain protocol such as the state channel, nodes which are not part of the state channel protocol do not know any details of the protocol’s execution due to privacy concerns. Onchain transactions on the hand are publicly accessible.
VIII Network Statistics
(a) Minimum accepted value of HTLC payment (msat)
(b) Distribution of base fee (msat)
(c) Distribution of rate fee (msat)
(d) Distribution of node degree
(e) Distribution of HTLC expiration time
Figure 5: $\mathsf{LN}$ statistics
This section introduces a general analysis of the parameters used in $\mathsf{LN}$, extracted from $\mathsf{LN}$ gossip messages [94]. It offers views into user behavior and the implications of such behavior for network performance. Our analysis shows that the nodes prefer to choose the default value of the $\mathsf{LN}$ parameters. Using standard default values for these parameters allows nodes to operate smoothly together, reducing the chance of compatibility issues and enhancing the overall reliability and efficiency of the $\mathsf{LN}$. We evaluated the analysis on a snapshot of $\mathsf{LN}$ topology from November 27, 2021 [95, 96]. The snapshot includes 18,331 nodes and 80,918 channels with a mean channel capacity of 4056152.35 satoshi and a median channel capacity is 1000000 satoshi (this provides us information on how funds are distributed among $\mathsf{LN}$ channels).
Although not explicitly shown in this chart, nodes with multiple connections suggest the presence of fewer central nodes compared to the end-users nodes. Figure 5(a) shows that the majority of the nodes (87.9%) set the ’htlc_minimum_msat’ to a default value of 1000 msat, which indicates that the $\mathsf{LN}$ supports micropayment transactions. Figure 5(b) shows that ’fee_base_msat’ parameter across all of the nodes has the highest two values 1000 msat as a default value, and 0 msat as the minimum value, which indicates that most of the nodes do not set the fee_base_msat beyond the default configuration, and likely to support the network usage. Figure 5(c) shows the majority of the nodes (51.8%) in $\mathsf{LN}$ use various ’fee_rate_milli_msat’ values. A substantial portion of nodes (35.2%) set their fee rate to 1 msat to attract more transactions. Also, different nodes have different fee rate strategies, making $\mathsf{LN}$ a competitive environment. Figure 5(d) shows that the majority of the nodes (49.5%) have only one connection that indicates that in $\mathsf{LN}$, many end-users connected to fewer well-connected nodes. Moreover, 30% of the nodes have no payment channels, these could be new nodes added to the $\mathsf{LN}$ or they could be inactive nodes. Figure 5(e) shows that (81.5%) of nodes prefer to set the ’cltv_expiry_delta’ to the default value of 40 blocks, which is the number of the blocks a node can wait before losing its satoshis from $\mathsf{LN}$. Moreover, it helps to increase the flexibility of the $\mathsf{LN}$ to include more nodes along the payment path before reaching the ”locktime max’ parameter value (2016 blocks). According to the analysis, the nodes generally prefer to use the default parameter values of the lightning network. Using these default values can help in developing more effective routing protocols and optimizing routing fee mechanisms for the nodes.
IX Conclusion
In this paper, we qualitatively compared the recent work in various areas of PCN research: pathfinding and routing, virtual channels, payment channel hubs, state channels and rebalancing. We have also provided the gaps in research in these areas along with reasons why fulfilling those gaps is a hard research challenge. We hope that this paper motivates researchers to build robust protocols that address these gaps that would go a long way towards building out and developing the decentralized finance ecosystem.